__17. A rectangular solid whose dimensions are 4cm, 8cm, and 12cm is inscribed in a sphere. What is the diagonal of the rectangular solid?

_18. In question 17, what is the surface area of the sphere in ¹?
1
Expert's answer
2020-12-16T08:11:30-0500

17. Let "a=" the length of a rectangular solid, "b=" its width, a "c=" its height.

The the diagonal "d" of a rectangular solid will be


"d=\sqrt{a^2 +b^2+c^2}""d=\sqrt{4^2+8^2+12^2}=4\sqrt{14}(cm)\approx14.97(cm)"



18. If the rectangular solid is inscribed in a sphere, the center of the sphere is the point of intersection of the diagonals of the rectangular solid. Then


"R_{sphere}=\dfrac{1}{2}d=\dfrac{1}{2}(4\sqrt{14}cm)=2\sqrt{14}\ cm"

The surface area of the sphere is


"A=4\pi R_{sphere}^2=4\pi (2\sqrt{14}cm)^2""=224\pi\ cm^2\approx703.72\ cm^2"